OpenStudy (anonymous):

sara is reading a 280-page book. if she has read 124 pages, what percent of the book has she not read

hero (hero):

\[\frac{280-124}{280} \times 100\]

OpenStudy (anonymous):

(Pages unread)/(total pages) *100

hero (hero):

This is similar to percent increase/decrease

hero (hero):

But in general it is a percent change problem with the general form: \[\% \Delta = \frac{|a - b|}{a} \times 100\] where: a = original amount b = new amount